# TC-KANRecon: High-Quality and Accelerated MRI Reconstruction via Adaptive KAN Mechanisms and Intelligent Feature Scaling

### v1

@@ 4.4 Comparison Experiments
Compared to the DiffuseRecon model, which achieves better reconstruction through the diffusion model, TC-KANRecon proves superior in handling complex textures. DiffuseRecon is prone to artifacts that lead to texture inconsistency and detail loss. For instance, the NMSE of DiffuseRecon is 0.0272, whereas TC-KANRecon achieves a lower NMSE of 0.0211, indicating higher image quality. The TC-KANRecon model combines a multi-attention mechanism with adaptive feature adjustment strategies, enhancing feature extraction and interpretability. This allows it to excel in capturing complex pathological features and restoring details. In contrast, the TC-DiffRecon model, although it addresses issues like image fragmentation caused by over-smoothing and improves overall image quality and consistency, still struggles with high-frequency details and complex textures. TC-KANRecon significantly improves the ability to capture complex pathological features and overall image reconstruction quality through its innovative Tok-KAN module and dynamic clipping strategy. Specifically, TC-KANRecon improves PSNR by 0.72 and SSIM by 0.076 on the STM-TEA dataset. DiffuseRecon’s NMSE is 0.0272, while TC-DiffRecon’s is 0.0247. In summary, TC-KANRecon excels in reproducing image details and handling complex textures, demonstrating higher robustness and stability. It maintains excellent performance across different acceleration factors, outperforming other models in various key metrics.
fastMRI [41] | |||||||||||||
Method | PSNR() | SSIM() | NMSE() | ||||||||||
AF Factor | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | |
ZF | 22.91 | 20.79 | 18.37 | 15.07 | 0.549 | 0.482 | 0.414 | 0.349 | 0.0436 | 0.0488 | 0.0567 | 0.0719 | |
KIKI-net [43] | 27.01 | 26.55 | 25.08 | 24.06 | 0.613 | 0.503 | 0.492 | 0.412 | 0.0354 | 0.0397 | 0.0462 | 0.0542 | |
D5C5 [16] | 27.74 | 26.56 | 26.12 | 24.60 | 0.632 | 0.529 | 0.512 | 0.445 | 0.0289 | 0.0374 | 0.0440 | 0.0532 | |
U-net [12] | 27.68 | 26.23 | 25.83 | 24.49 | 0.631 | 0.520 | 0.511 | 0.436 | 0.0295 | 0.0382 | 0.0430 | 0.0520 | |
OUCR [22] | 28.08 | 27.66 | 26.23 | 25.31 | 0.633 | 0.534 | 0.516 | 0.482 | 0.0277 | 0.0368 | 0.0389 | 0.0486 | |
DiffuseRecon [32] | 28.27 | 27.83 | 26.60 | 25.90 | 0.635 | 0.586 | 0.549 | 0.528 | 0.0272 | 0.0360 | 0.0381 | 0.0466 | |
ReconFormer [44] | 28.62 | 28.01 | 26.64 | 26.08 | 0.643 | 0.590 | 0.529 | 0.519 | 0.0263 | 0.0348 | 0.0374 | 0.0480 | |
TC-DiffRecon [45] | 29.78 | 28.77 | 27.35 | 26.49 | 0.742 | 0.637 | 0.591 | 0.562 | 0.0247 | 0.0313 | 0.0338 | 0.0415 | |
SKM-TEA [42] | |||||||||||||
Method | PSNR() | SSIM() | NMSE() | ||||||||||
AF Factor | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | |
ZF | 23.63 | 22.09 | 20.06 | 16.72 | 0.653 | 0.587 | 0.529 | 0455 | 0.0361 | 0.0422 | 0.0511 | 0.0641 | |
KIKI-net [43] | 28.16 | 26.96 | 25.80 | 25.11 | 0.716 | 0.620 | 0.601 | 0.543 | 0.0286 | 0.0350 | 0.0434 | 0.0510 | |
D5C5 [16] | 28.87 | 28.24 | 27.35 | 27.01 | 0.731 | 0.654 | 0.637 | 0.586 | 0.0221 | 0.0260 | 0.0384 | 0.0460 | |
U-net [12] | 28.52 | 27.79 | 26.73 | 26.34 | 0.723 | 0.645 | 0.631 | 0.571 | 0.0254 | 0.0336 | 0.0410 | 0.0493 | |
OUCR [22] | 29.26 | 28.90 | 28.06 | 27.24 | 0.745 | 0.656 | 0.642 | 0.603 | 0.0165 | 0.0247 | 0.0320 | 0.0387 | |
DiffuseRecon [32] | 29.51 | 29.13 | 28.20 | 27.49 | 0.749 | 0.712 | 0.682 | 0.668 | 0.0152 | 0.0232 | 0.0303 | 0.0352 | |
ReconFormer [44] | 29.74 | 29.43 | 28.77 | 28.37 | 0.765 | 0.724 | 0.693 | 0.654 | 0.0149 | 0.0216 | 0.0278 | 0.0326 | |
TC-DiffRecon [45] | 30.59 | 30.26 | 29.50 | 29.22 | 0.852 | 0.771 | 0.732 | 0.708 | 0.0136 | 0.0198 | 0.0254 | 0.0311 | |
@@ 4.5 Validation of Model Generalizability
fastMRI | SKM-TEA | ||||||||||||
@@
4.6 Ablation Experiments
4.6.1 Impact after removing the MF Model
@@ 5 Conclusion
Given the existing challenges in MR images analysis, particularly in the efficient processing and accurate reconstruction of knee MRI datasets, this paper presents an innovative deep learning architecture: TC-KANRecon. This architecture addresses the limitations of traditional MRI reconstruction models, such as inadequate feature extraction and poor robustness. The TC-KANRecon framework incorporates multi-head attention mechanisms and scalar modulation factors to finely control backbone and skip features, effectively balancing denoising needs while preserving structural information. This enhances the model’s robustness and retention of structure in complex noise environments. Additionally, the MC-Model module within the TC-KANRecon framework integrates fully sampled k-space information with MRI data, further improving the realism and richness of reconstructed images. The proposed dynamic clipping strategy adjusts the cropping interval based on sampling steps, mitigating image detail loss caused by traditional cropping methods and enriching the visual features of reconstructed images.
To validate the effectiveness of the TC-KANRecon model, this study employs two large-scale knee MRI datasets: fastMRI, comprising 1,172 subjects and over 41,020 slice data, and the SKM-TEA dataset, containing 155 subjects and over 24,800 slice data. The datasets were rigorously split into training and evaluation sets to ensure the objectivity and reliability of the experimental results. The findings demonstrate that the TC-KANRecon model outperforms existing methods in reconstructed image quality, achieving an optimal balance between denoising and detail preservation. In addition, we also further confirmed the critical role of each component in the overall performance of the model through ablation experiments, particularly highlighting the significant contributions of multi-head attention mechanisms and scalar modulation factors in enhancing feature extraction and denoising capabilities.
The TC-KANRecon model provides a novel approach for efficiently processing and accurately reconstructing MR images. Despite these significant advancements, there are still some limitations and areas for improvement. One major challenge is the model’s reliance on large-scale datasets for training, which can be problematic in situations where such extensive data is unavailable. Future work will explore techniques such as data augmentation, transfer learning, and synthetic data generation to mitigate this issue. Another important consideration is the computational efficiency of the TC-KANRecon model. As deep learning models grow increasingly complex, it becomes crucial to ensure they can be deployed in real-time clinical environments without significant delays. Future research will focus on optimizing the model architecture to accelerate inference time while maintaining high reconstruction quality.

### v2

@@ 4.4 Comparison Experiments
Compared to the DiffuseRecon model [35], which achieves better reconstruction through the diffusion model, TC-KANRecon proves superior in handling complex textures. DiffuseRecon is prone to artifacts that lead to texture inconsistency and detail loss. For instance, the NMSE of DiffuseRecon is 0.0272, whereas TC-KANRecon achieves a lower NMSE of 0.0211, indicating higher image quality. The TC-KANRecon model combines a multi-attention mechanism with adaptive feature adjustment strategies, enhancing feature extraction and interpretability. This allows it to excel in capturing complex pathological features and restoring details. In contrast, the TC-DiffRecon model, although it addresses issues like image fragmentation caused by over-smoothing and improves overall image quality and consistency, still struggles with high-frequency details and complex textures. TC-KANRecon significantly improves the ability to capture complex pathological features and overall image reconstruction quality through its innovative Tok-KAN module and dynamic clipping strategy. Specifically, TC-KANRecon improves PSNR by 0.72 and SSIM by 0.076 on the STM-TEA dataset. DiffuseRecon’s NMSE is 0.0272, while TC-DiffRecon’s is 0.0247. In summary, TC-KANRecon excels in reproducing image details and handling complex textures, demonstrating higher robustness and stability. It maintains excellent performance across different acceleration factors, outperforming other models in various key metrics.
@@ 4.5 Validation of Model Generalizability
4.6 Ablation Study
4.6.1 Impact after removing the MF Model
fastMRI [45] | SKM-TEA [46] | ||||||||||||
@@ 5 Conclusion
This paper introduces an innovative deep learning framework, TC-KANRecon, designed to accelerate and enhance the quality of MRI image reconstruction. By integrating the MF-UKAN module and a dynamic clipping strategy, the framework effectively balances denoising and structural preservation, significantly improving the model’s robustness and retention of structural details in complex noise environments. Additionally, the MC-Model module incorporates fully sampled k-space information, further enhancing the realism and richness of reconstructed images. Experimental results on two large-scale knee MRI datasets, fastMRI and SKM-TEA, demonstrate that TC-KANRecon outperforms existing methods in terms of image quality, achieving an optimal balance between denoising and detail preservation, and showcasing strong generalization capabilities. The significance of this work lies in providing a highly efficient and accurate MRI image reconstruction method for clinical diagnosis, potentially optimizing the use of medical resources and enhancing diagnostic accuracy.
Despite these significant advancements, there are still some limitations and areas for improvement. One major challenge is the model’s reliance on large-scale datasets for training, which can be problematic in situations where such extensive data is unavailable. Future work will explore techniques such as data augmentation, transfer learning, and synthetic data generation to mitigate this issue. Another important consideration is the computational efficiency of the TC-KANRecon model. As deep learning models grow increasingly complex, it becomes crucial to ensure they can be deployed in real-time clinical environments without significant delays. Future research will focus on optimizing the model architecture to accelerate inference time while maintaining high quality.